# The Monge optimal transport barycenter problem

### v1

@@ 8 Conditional density estimation
Our methodology simulates $\rho(x|z)$, by producing $n$ samples $\{x^{*}_{i}\}$ from $\rho(x|z^{*})$ for any target $z^{*}$. Simulation is at the core of many applications, but others, such as Bayesian inference, require the evaluation of $\rho(x|z)$ for arbitrary values of $x$ and $z$. The fact that typically there is none or at most one observation available for any target value $z$ makes estimating $\rho(x|z)$ directly from the data $\{x_{i},z_{i}\}$ challenging. A slight extension of our procedure produces such conditional density estimation.
Regular –as opposed to conditional– density estimation can be obtained through optimal transport as follows. Given $n$ samples $\{x_{i}\}$ drawn from the unknown distribution $\rho(x)$ that we seek to estimate, select a target distribution $\mu(y)$ that one can easily both evaluate and sample, such as a Gaussian, and find the optimal map $Q(x)$ pushing forward $\rho$ to $\mu$. Then
@@
If the density $\rho(x)$ is sought for values of $x$ different from the $\{x_{i}\}$, one can carry these values through the procedure to their final $y=Q(x)$ as passive tracers that do not affect $L$ in (18). Alternatively, one can solve the reciprocal optimal transport problem from $\mu$ to $\rho$, and then write
$\rho(x)=\left|\det\left(\nabla_{x}Y(x)\right)\right|\ \mu(Y(x))$
for any value of $x$ sought.
This density estimation procedure requires selecting a target measure $\mu=\rho_{1}$. We can either adopt a fixed target, such as a standard Gaussian, or to adapt it to the data, using for instance a Gaussian with the same mean and covariance matrix as the data or a Gaussian mixture fitted to the data through Expectation Maximization. The advantage of such more tailored approaches is that the corresponding optimal transport problem becomes easier, since even the trivial map $Q(x)=x$ provides a regular parametric density estimation.
@@
    Obtain $n$ samples $\{x^{*}_{i}\}$ from $\rho(x|z^{*})$ and apply density estimation to these directly.
@@
One would choose the first approach when seeking $\rho(x|z)$ for many values of $x$ and only a handful of values of $z$, and the second approach when exploring the dependence of the conditional density on $z$, as beholds for instance Bayesian inference.
@@ 9 Numerical examples
We illustrate the methodology through numerical examples, using synthetic data first and then two real data applications: uncovering hidden patterns in the atmospheric temperature at ground level and forecasting ocean states.
@@ 9.1 A Gaussian distribution with $z$-dependent mean and variance
with $z=\left(z_{1},z_{2}\right)$ uniformly distributed in the square $-\frac{1}{2}\leq z_{1,2}\leq\frac{1}{2}$. The data is displayed on the top left panel of figure 3. Since for each value of $z$ the distribution for $x$ is Gaussian, it can be fully captured using the two-dimensional test function space $\mathcal{G}(y)$ spanned by the functions $y$ and $y^{2}$, while keeping for $\mathcal{F}(z)$ a general adaptive space based on kernels. We display the results of the run in figure 3 through the corresponding $\{y_{i}\}$, whose $z$-independent distribution is a Gaussian, as beholds the barycenter of a set of Gaussians, and the simulation and estimation of $\rho(x|z_{*})$ for two selected values of $z_{*}$, with the true underlying distribution also drawn for comparison.
Figure 3: First Example. The leftmost column displays the samples $\{x_{i}\}$ in the top row and the corresponding $\{y_{i}\}$ in the barycenter in the bottom row. The middle and rightmost columns show the simulated samples $\{x_{i}^{*}\}$ and the estimated versus the true density $\rho(x|z_{*})$ for $z_{*}=(0.2,0.3)$ and $z_{*}=(-0.2,-0.3)$.
@@ 9.2 Two $z$-dependent Gaussian mixtures
In order to consider non-Gaussian examples –more precisely, examples where the dependence of $x$ on $z$ cannot be reduced to a $z$-dependent translation and scaling– we perform first a one-dimensional experiment, drawing 1500 samples from the $z$-dependent Gaussian mixture
@@
The datapoints are displayed on the top left panel of figure 4.
This example requires a test function space $\mathcal{G}(y)$ that goes beyond linear and quadratic functions. We performed four successive barycenter problems, the first with just linear and quadratic $\mathcal{G}(y)$, the rest with adaptive kernels, with the bandwidths of the kernels for both $\mathcal{G}(y)$ and $\mathcal{F}(z)$ adopted smaller for each successive run. The results, displayed in figure 4, show how the simulated samples and conditional density estimation recover the original $z$-dependent Gaussian mixture.
Figure 4: One-dimensional Gaussian mixture. The leftmost column displays the data samples in the top row and the barycenter in the bottom row. The middle and rightmost columns show the simulated samples and the estimated versus the true density for $z_{*}=0.5$ and $z_{*}=-0.5$.
We extend this example to the two-dimensional Gaussian mixture
@@
The data points and the barycenter are visualized in figure 5 and the results of the procedure are displayed in figure 6.
Figure 5: A Two-dimensional Gaussian mixture. The left panel displays the data points, while the right panel shows the equivalent plots for the barycenter.
Figure 6: Two-dimensional Gaussian mixture. The top and bottom rows show the simulated samples, the estimated conditional density, and the true density for $z_{*}=0.5$ and $z_{*}=-1$ respectively.
@@ 9.3 An example of Bayesian inference
This section illustrates model-free Bayesian inference using the OTBP. A different use of push-forward maps for Bayesian inference [37] pushes forward the prior to the posterior measure. To demonstrate our approach, we draw samples from the distribution
$\rho(x|z)=N(z^{2}-2,\sigma^{2}),\quad\sigma=0.5,\quad z\sim\text{U}([-2,2]).$
The left panel of figure 7 displays the data $\{x_{i}\}$ and the discovered $\{y_{i}\}$ as functions of the corresponding $\{z_{i}\}$. From these, we can directly infer the distribution of $z$ given an observation $x$:
@@
where $\rho(x|z)$ is the not the distribution we know in closed form but the one inferred from the data through the OTBP. We have adopted as a natural default prior $\gamma_{pr}$ the distribution underlying the observed $\{z_{i}\}$. The results for two values of $x$ are displayed on the middle and right panels of Figure 7, overlapped with the exact answer. They succeed in capturing the transition from unimodal to bimodal distributions corresponding to the parabolic dependence of the conditional mean of $x$ on $z$.
@@
[…]
@@
where $\alpha\in(0,1)$ is an unknown model parameter and $\beta=\sigma=0.5$ are fixed drift and noise levels. Our goal is to learn the model from a set of training data pairs $(X_{\text{train}}^{n+1};\,X_{\text{train}}^{n},\alpha_{\text{train}}^{n})$ and use the model learned to estimate $\alpha$ online from a testing series, while making increasingly more accurate forecasts. For the training data, we draw $\alpha_{\text{train}}^{n}$ from a beta distribution $B(2,2)$ over $(0,1)$, which we also adopt as prior $\gamma^{0}(\alpha)$, and $X_{\text{train}}^{n}$ from the uniform distribution $\text{U}([a,b])$.
We carry out experiments with two different parameter values, $\alpha=0.2,0.8$, with the corresponding test data displayed in Figure 8. The results in Figure 9 demonstrate that the posterior densities converge to delta functions centered around the corresponding true parameters.
Figure 8: Time-discretized Ornstein–Uhlenbeck process. Testing time series for $\alpha=0.2,0.8$.
Figure 9: Estimated posterior densities as the number of time steps grows for $\alpha=0.2$ in the top row and $\alpha=0.8$ in the bottom row.
@@
$\frac{dx_{2}}{dt}$ $=-\gamma x+\delta x_{1}x_{2},$
@@
We adopt as test data a simulation with parameters $\alpha=0.3,\beta=0.9,\gamma=0.5,\delta=0.4$ and $\epsilon=0.1$, yielding the periodic results displayed in Figure 10. For a first experiment, we take $\alpha$ as the only unknown parameter. Again, random training data pairs $(X_{\text{train}}^{n+1};\,X_{\text{train}}^{n},\alpha_{\text{train}}^{n},\Delta t^{n})$ are generated, drawing $X_{\text{train}}^{n}$ from a uniform distribution, $\alpha_{\text{train}}^{n}$ from the beta distribution $B(2,2)$, and deriving the corresponding $X_{\text{train}}^{n+1}$ from the model with additive noise of level $\epsilon$.
As before, we learn the conditional density $\rho(X^{n+1}|X^{n},\alpha,\Delta t^{n})$ by solving the barycenter problem for the training data. Then we apply Bayes rule online to the testing data, updating at each step
@@
The results are shown in the top panel of Figure 11.
We consider next a situation where two parameters, $\alpha$ and $\gamma$, are unknown, so the training data consists of quintuples $(X_{\text{train}}^{n+1};\,X_{\text{train}}^{n},\alpha_{\text{train}}^{n},\gamma_{\text{train}}^{n},\Delta t^{n})$, and the joint posterior density should be estimated through
$\rho^{n+1}(\alpha,\gamma)\propto\rho(X^{n+1}|X^{n},\alpha,\gamma,\Delta t^{n})\cdot\rho^{n}(\alpha,\gamma),$
with joint Gaussian prior $\rho^{0}(\alpha,\gamma)=N\left(\begin{pmatrix}0.5\\ 0.5\end{pmatrix},\begin{pmatrix}0.2&0\\ 0&0.2\end{pmatrix}\right)$. The results are displayed in the bottom panel of Figure 11. We see how in both cases the estimation converges to a delta function centered at the right underlying value of the parameters.
Figure 10: Lotka-Volterra model: Testing time series for $\alpha=0.3,\beta=0.9,\gamma=0.5,\delta=0.4$.
Figure 11: Estimated posterior densities as the number of time steps grows for the single unknown $\alpha=0.3$ in the top row, and for two unknowns $\alpha=0.3,\gamma=0.5$ in the bottom row.
@@ 9.5 Uncovering a hidden signal
Recall from Theorem 3 that the solution $y=T(x,z)$ of the barycenter problem is a proxy for the variable $w$, with $\phi(z,w)\rightarrow X(y,z)$. Moreover, $y$ is related to any “true” hidden variable $w$ through a possibly $z$-dependent function
@@
Figure 12 presents three synthetic examples in order to illustrate the different kind of dependence between $y$ and $w$ typically observed in applications. The panels of each row are relative to the different synthetic examples described below. The panels on the left column display the $\{x_{i}\}$ and corresponding $\{y_{i}\}$ in terms of the $\{z_{i}\}$, the middle column displays $y(z)$ again, colored according to the corresponding value of $w$, and the right column displays $y(w)$, colored according to $z$. In the first row, $z\sim\text{U}[0.25,1]$, $w\sim\text{U}[-1,1]$ and $x=\phi(z,w)=zw^{3}$ (we exclude values of $z$ near 00 because $\forall w\ \phi(0,w)=0$, i.e. $\rho(x|0)$ does not vanish on small sets.) In this example $Y_{z}(w)=Y(w)$ does not depend on $z$ and $Y(w)$ is invertible. In the second row $x=zw^{2}$ under the same distributions for $z$ and $w$. We still have that $Y_{z}(w)$ does not depend on $z$ but now $Y(w)$ is not globally invertible, a reflection of the fact that the sign of $w$ is not identifiable, since $\forall z$ and $\forall w$ we have that $\phi(z,-w)=\phi(z,w)$. In the third row, $w\sim\text{U}[0,1]$, $z\sim\text{U}[(-1,-0.25)\cup(0.25,1)]$ and $x=zw$, for which $Y_{z}(w)$ depends on the sign of $z$ (we leave it as a challenge to the interested reader to uncover why this is so.)
@@
The analysis of the barycenter underlying the points $\{y_{i}\}$ may at first seem similar to residual analysis, whereby the difference between actual and predicted values is further analyzed to assess model adequacy and improve its predictive power [38, 39]. Both procedures aim to remove variability in the data $x$ attributed to the cofactors $z$, yet while residual analysis only removes $\bar{x}(z)$, the conditional expected value of $x$, the barycenter does this at the level of the full probability distributions $\rho(x|z)$ underlying the data. Consider the example in the first row of Figure 12, where the data are generated according to $x=zw^{3}$ with $z$ and $w$ drawn uniformly. The distribution of the residuals (obtained by subtracting the regression of $x$ vs $z$ from the actual value of $x$) would, in this case, be identical to the original distribution underlaying the $x$, providing no new useful information. By contrast, the points $y_{i}$ in the barycenter represent the full variability of $x$ not explainable by $z$. If the $y$’s can be related to known factors $w$, this can be used to improve the model for $x$, for instance by regressing $y$ against these factors and then using both $z$ and the reconstructed $y$ to predict $x$. Better still, instead of regression, the barycenter problem can be used once again to simulate $\rho(y|w)$.

### v2

@@ 8 Conditional density estimation
Our methodology simulates $\rho(x|z)$, by producing $n$ samples $\{x^{*}_{i}\}$ from $\rho(x|z_{*})$ for any target $z_{*}$. Simulation is at the core of many applications, but others, such as Bayesian inference, require the evaluation of $\rho(x|z)$ for arbitrary values of $x$ and $z$. The fact that typically there is none or at most one observation available for any target value $z$ makes estimating $\rho(x|z)$ directly from the data $\{x_{i},z_{i}\}$ challenging. A slight extension of our procedure produces such conditional density estimation.
Regular –as opposed to conditional– density estimation can be obtained through the OTP as follows. Given $n$ samples $\{x_{i}\}$ drawn from the unknown distribution $\rho(x)$ that we seek to estimate, select a target distribution $\mu(y)$ that one can easily both evaluate and sample, such as a Gaussian, and find the optimal map $Q(x)$ pushing forward $\rho$ to $\mu$. Then
@@
If the density $\rho(x)$ is sought for values of $x$ different from the $\{x_{i}\}$, one can carry these values through the procedure to their final $y=Q(x)$ as passive tracers that do not affect $L$ in (17). Alternatively, one can solve the reciprocal OTP from $\mu$ to $\rho$, and then write $\rho(x)=\left|\det\left(\nabla_{x}Y(x)\right)\right|\ \mu(Y(x))$ for any value of $x$ sought.
This density estimation procedure requires selecting a target measure $\mu=\rho_{1}$. We can either adopt a fixed target, such as a standard Gaussian, or adapt it to the data, using for instance a Gaussian with the same mean and covariance matrix as the data or a Gaussian mixture fitted to the data through Expectation Maximization. The advantage of such more tailored approaches is that the corresponding OTP becomes easier, since even the trivial map $Q(x)=x$ provides a regular parametric density estimation.
@@
    Obtain $n$ samples $\{x^{*}_{i}\}$ from $\rho(x|z_{*})$ and apply density estimation to these directly.
@@
One would choose the first approach when seeking $\rho(x|z)$ for many values of $x$ and only a handful of values of $z$, and the second when exploring the dependence of the conditional density on $z$, as in Bayesian inference.
@@ 9 Numerical examples
We illustrate the methodology through numerical examples, using both synthetic and real data.
@@ 9.1 A Gaussian distribution with $z$-dependent mean and variance
with $z=\left(z_{1},z_{2}\right)$ uniformly distributed in the square $\frac{1}{2}\leq z_{1,2}\leq\frac{1}{2}$, displayed on the top left panel of figure 3(a). Since for each value of $z$ the distribution for $x$ is Gaussian, it can be fully captured using the two-dimensional test function space $\mathcal{G}(y)$ spanned by the functions $y$ and $y^{2}$, while keeping for $\mathcal{F}(z)$ a general adaptive space based on kernels. We display the results of the run in figure 3(a) through the corresponding $\{y_{i}\}$ and the simulation and estimation of $\rho(x|z_{*})$ for two selected values of $z_{*}$, with the true underlying distribution also drawn for comparison.
(a) First Example. The leftmost column displays the samples $\{x_{i}\}$ in the top row and the $\{y_{i}\}$ in the barycenter in the bottom row, with corresponding histograms showing that the final barycenter converges to a Gaussian. The middle and rightmost columns show the simulated samples $\{x_{i}^{*}\}$ and the estimated versus the true density $\rho(x|z_{*})$ for $z_{*}=(0.2,0.3)$ and $z_{*}=(-0.2,-0.3)$.
(b) One-dimensional Gaussian mixture. The leftmost column displays the data samples in the top row and the barycenter in the bottom row. The middle and rightmost columns show the simulated samples and the estimated versus the true density for $z_{*}=0.5$ and $z_{*}=-0.5$.
Figure 3: 1D Examples.
@@ 9.2 Two $z$-dependent Gaussian mixtures
In order to consider non-Gaussian examples –more generally, examples where the dependence of $x$ on $z$ cannot be reduced to a $z$-dependent linear transformation– we perform first a one-dimensional experiment, drawing 1500 samples from the $z$-dependent Gaussian mixture, displayed on the top left panel of figure 3(b):
@@
This example requires a test function space $\mathcal{G}(y)$ that goes beyond linear and quadratic functions. We performed four successive barycenter problems, the first with just linear and quadratic $\mathcal{G}(y)$, the rest with adaptive kernels, with the bandwidths of the kernels for both $\mathcal{G}(y)$ and $\mathcal{F}(z)$ adopted smaller for each successive run. The results, displayed in figure 3(b), show how the simulated samples and conditional density estimation recover the original $z$-dependent Gaussian mixture.
We extend this example to the two-dimensional Gaussian mixture, visualized in panel (a) of figure 4:
@@
The results of the procedure are displayed in panel (b) of figure 4 .
(a) A two-dimensional Gaussian mixture. The left panel displays the data points, while the right panel shows the equivalent plots for the barycenter.
(b) Two-dimensional Gaussian mixture. The top and bottom rows show the simulated samples, the estimated conditional density, and the true density for $z_{*}=0.5$ and $z_{*}=-1$, respectively.
Figure 4: Two-dimensional Gaussian mixture.
@@ 9.3 An example of Bayesian inference
This section illustrates model-free Bayesian inference using the OTBP (A different use of push-forward maps for Bayesian inference [38] pushes forward the prior to the posterior measure.) To demonstrate our approach, we draw samples from the distribution $\rho(x|z)=N(z^{2}-2,\sigma^{2})$, $\sigma=0.5$, $z\sim\text{U}([-2,2])$. The left panel of figure 5 displays the data $\{x_{i}\}$ and the discovered $\{y_{i}\}$ as functions of the corresponding $\{z_{i}\}$. From these, we can directly infer the distribution of $z$ given an observation $x$:
@@
where $\rho(x|z)$ is the distribution inferred from the data through the OTBP. We have adopted as prior $\gamma_{pr}$ the distribution underlying the observed $\{z_{i}\}$. The results for two values of $x$ are displayed on the middle and right panels of Figure 5, overlapped with the exact answer. They succeed in capturing the transition from unimodal to bimodal distributions corresponding to the parabolic dependence of the conditional mean of $x$ on $z$.
@@
with the proportionality constant determined by the normalizing condition that $\int d\gamma(\alpha)=1$. In the conventional setting, the conditional distribution $\rho$ is known except for the parameters $\alpha$. We can extend this framework to situations where $\rho(X^{n+1}|X^{n},Z^{n},\alpha)$ itself is only known from partial observations of time series under different values of $\alpha$ and $Z$. In a medical setting for instance, $X$ may represent glucose concentration in the bloodstream, $Z^{n}$ the caloric intake at time $t_{n}$, and $\alpha$ a patient’s parameter that may only be determined after treatment. Having observed in the past a number of patients under different diets and having determined their corresponding parameters $\alpha$, we can use this for the online estimation of $\alpha$ for a patient currently under treatment.
@@
where $\alpha\in(0,1)$ is an unknown model parameter and $\beta=\sigma=0.5$ are fixed drift and noise levels. Our goal is to learn the model from a set of training data pairs $(X_{\text{train}}^{n+1};\,X_{\text{train}}^{n},\alpha_{\text{train}}^{n})$ and use the model learned to estimate $\alpha$ online from a testing series, while making increasingly more accurate forecasts. For the training data, we draw $\alpha_{\text{train}}^{n}$ from a beta distribution $B(2,2)$ over $(0,1)$, which we also adopt as prior $\gamma^{0}(\alpha)$, and $X_{\text{train}}^{n}$ from the uniform distribution $\text{U}([a,b])$. We carry out experiments with two different parameter values, $\alpha=0.2,0.8$, with the corresponding test data displayed in the leftmost column in Figure 6. The following results in Figure 6 demonstrate that the posterior densities converge to delta functions centered around the corresponding true parameters.
Figure 6: Time-discretized Ornstein–Uhlenbeck process. Testing time series for $\alpha=0.2,0.8$. Estimated posterior densities as the number of time steps grows for $\alpha=0.2$ in the top row and $\alpha=0.8$ in the bottom row.
@@
$\frac{dx_{2}}{dt}$ $=-\gamma x_{2}+\delta x_{1}x_{2},$
@@
We adopt as test data a simulation with parameters $\alpha=0.3,\beta=0.9,\gamma=0.5,\delta=0.4$ and $\epsilon=0.1$, yielding the periodic results displayed in the leftmost panel of Figure 7. For a first experiment, we take $\alpha$ as the only unknown parameter. Again, random training data pairs $(X_{\text{train}}^{n+1};\,X_{\text{train}}^{n},\alpha_{\text{train}}^{n},\Delta t^{n})$ are generated, drawing $X_{\text{train}}^{n}$ from a uniform distribution, $\alpha_{\text{train}}^{n}$ from the beta distribution $B(2,2)$, and deriving the corresponding $X_{\text{train}}^{n+1}$ from the model with additive noise of level $\epsilon$. As before, we learn the conditional density $\rho(X^{n+1}|X^{n},\alpha,\Delta t^{n})$ by solving the barycenter problem for the training data. Then we apply Bayes rule online to the testing data, updating at each step
@@
The results are shown in the top row of Figure 7. We consider next a situation where two parameters, $\alpha$ and $\gamma$, are unknown, so the training data consists of quintuples $(X_{\text{train}}^{n+1};\,X_{\text{train}}^{n},\alpha_{\text{train}}^{n},\gamma_{\text{train}}^{n},\Delta t^{n})$, and the joint posterior density should be estimated through $\rho^{n+1}(\alpha,\gamma)\propto\rho(X^{n+1}|X^{n},\alpha,\gamma,\Delta t^{n})\cdot\rho^{n}(\alpha,\gamma)$, with joint Gaussian prior $\rho^{0}(\alpha,\gamma)=N\left(\begin{pmatrix}0.5\\ 0.5\end{pmatrix},\begin{pmatrix}0.2&0\\ 0&0.2\end{pmatrix}\right)$. The results are displayed in the bottom row of Figure 7. We see how in both cases the estimation converges to a delta function centered at the right underlying value of the parameters.
Figure 7: Lotka-Volterra model: Testing time series for $\alpha=0.3,\beta=0.9,\gamma=0.5,\delta=0.4$. Estimated posterior densities as the number of time steps grows for the single unknown $\alpha=0.3$ in the top row, and for two unknowns $\alpha=0.3,\gamma=0.5$ in the bottom row.
@@ 9.5 Uncovering a hidden signal
Recall from Lemma 2 that the solution $y=T(x,z)$ of the barycenter problem is a proxy for the variable $w$, with $\phi(z,w)\rightarrow X(y,z)$. Moreover, $y$ is related to any “true” hidden variable $w$ through a possibly $z$-dependent function
@@
Figure 8 presents three synthetic examples in order to illustrate the different kind of dependence between $y$ and $w$ typically observed in applications. The panels of each row correspond to the different synthetic examples described below. The left column display the $\{x_{i}\}$ and corresponding $\{y_{i}\}$ in terms of the $\{z_{i}\}$, the middle column displays $y(z)$ again, colored according to the corresponding value of $w$, and the right column displays $y(w)$, colored according to $z$. In the first row, $z\sim\text{U}[0.25,1]$, $w\sim\text{U}[-1,1]$ and $x=\phi(z,w)=zw^{3}$ (we exclude values of $z$ near 0 because $\forall w\ \phi(0,w)=0$, i.e. $\rho(x|0)$ does not vanish on small sets.) In this example $Y_{z}(w)=Y(w)$ does not depend on $z$ and $Y(w)$ is invertible. In the second row $x=zw^{2}$ under the same distributions for $z$ and $w$. We still have that $Y_{z}(w)$ does not depend on $z$ but now $Y(w)$ is not globally invertible, a reflection of the fact that the sign of $w$ is not identifiable, since $\forall z$ and $\forall w$ we have that $\phi(z,-w)=\phi(z,w)$. In the third row, $w\sim\text{U}[0,1]$, $z\sim\text{U}[(-1,-0.25)\cup(0.25,1)]$ and $x=zw$, for which $Y_{z}(w)$ depends on the sign of $z$.
@@
The analysis of the barycenter underlying the points $\{y_{i}\}$ may at first seem similar to residual analysis, whereby the difference between actual and predicted values is further analyzed to assess model adequacy and improve its predictive power [39, 40]. Both procedures aim to remove variability in the data $x$ attributed to the cofactors $z$, yet while residual analysis only removes $\bar{x}(z)$, the conditional expected value of $x$, the barycenter does this at the level of the full probability distributions $\rho(x|z)$ underlying the data. Consider the example in the first row of Figure 8, where the data are generated according to $x=zw^{3}$ with $z$ and $w$ drawn uniformly. The distribution of the residuals (obtained by subtracting the regression of $x$ vs $z$ from the actual value of $x$) would, in this case, be identical to the original distribution underlying the $x$, providing no new useful information. By contrast, the points $y_{i}$ in the barycenter represent the full variability of $x$ not explainable by $z$. If the $y$’s can be related to known factors $w$, this can be used to improve the model for $x$, for instance by regressing $y$ against these factors and then using both $z$ and the reconstructed $y$ to predict $x$. Better still, instead of regression, the barycenter problem can be used once again to simulate $\rho(y|w)$.